# A non-abelian version of Deligne’s Fixed Part Theorem

### v4

@@ Abstract.
We formulate and prove a non-abelian analog of Deligne’s Fixed Part theorem on Hodge classes, revisiting previous work of Jost–Zuo, Katzarkov–Pantev and Landesman–Litt. To this aim we study algebraically isomonodromic extensions of local systems and we relate them to variations of Hodge structures, for example we show that the Mumford-Tate group at a generic point stays constant in an algebraically isomonodromic extension of a variation of Hodge structure.
@@ 1. Introduction
On a quasi-projective complex manifold, a rational Betti class is called Hodge if it is integral and lies as a de Rham class in the right levels of the Hodge and the weight filtrations. Let $f\colon X\to S$ be a projective morphism between quasi-projective complex manifolds. Deligne’s Fixed Part Theorem asserts that a Hodge class on a fibre $X_{s}$ has finite orbit under the fundamental group of $S$ based at $s$ if and only if it extends to a Hodge class on $X$ after base change to a finite étale cover of $S$. Here we have to make the technical assumption that $S$ topologically has the properties of Artin’s elementary neighborboods, see Theorem 3.1 for a precise formulation.
@@
If $R\subset K$ is a subring we denote by $(R)K\text{-}{\sf PVHS}$ those polarized variations of $K$-Hodge structure $(\mathbb{L},{\mathsf{F}},Q)$ such that the local system $\mathbb{L}$ can be defined over $R$.
Let $\overline{f}\colon\overline{X}\to S$ be a projective holomorphic submersion of connected complex manifolds with connected fibres. Let $X\subset\overline{X}$ be the complement of a relative simple normal crossings divisor $D\subset\overline{X}$ over $S$. We call $f=\overline{f}|_{X}$ a good holomorphic map. Throughout the introduction we assume that $S$ is quasi-projective and that one of the following properties holds for $f$.
@@
Let $f:X\to S$ be a good holomorphic map, $s$ be a point in $S$. Let $(\mathbb{L}_{s},{\mathsf{F}}_{s},Q_{s})$ be a $({\mathbb{Z}})K\text{-}{\sf PVHS}$ on the fibre $X_{s}$, where $K\in\{{\mathbb{Q}},{\mathbb{R}},{\mathbb{C}}\}$. Then the following conditions are equivalent:
@@
    the filtration ${\mathsf{F}}_{s}$ extends to a (relative) Griffiths transversal filtration by subbundles $\hat{\mathsf{F}}$ of the formal isomonodromic Deligne extension (^,^:^$(\hat{\mathcal{E}},\hat{\nabla}\colon\hat{\mathcal{E}}\to\hat{\Omega}^{1}_{\hat{\overline{X}}/\hat{S}}(\log D)(\hat{\mathcal{E}}))$.
Here “essentially extends” means that it extends after base change with respect to a finite, étale, surjective covering of $S$, see Subsection 3.1 for a precise formulation. In 5) $\hat{S}$ is the formal scheme of $S$ along $s$ and $\hat{\overline{X}}$ the formal scheme of $\overline{X}$ along $\overline{X}_{s}$. The formal isomonodromic Deligne extension $(\hat{\mathcal{E}},\hat{\nabla})$ is the formal completion of the log flat bundle on $\overline{X}_{\Delta}$ associated by Deligne [Del70, Prop. 5.4, Rmk. 5.5] to the local system on $X_{\Delta}$ which canonically extends $\mathbb{L}_{s}$. Recall that the filtration ${\mathsf{F}}_{s}$ canonically extends to a filtration by subbundles of the Deligne extension over $\overline{X}_{s}$ by Schmid’s nilpotent orbit theorem, see [CK89, (2.1)], so part 5) can be interpreted as a formal deformation problem for the latter filtration of bundles from $\overline{X}_{s}$ to $\hat{\overline{X}}$.
@@
In order to show the implication 4) $\Rightarrow$ 1) we use Simpson’s non-abelian Hodge loci. In fact he considers only the case $f$ projective and we generalize his theory to the case of a good holomorphic map $f$ in Appendix A. The analogous implication 5) $\Rightarrow$ 1) was suggested to us by D. Litt, see also [Lit24, Cor. 4.4.3].
@@
As mentioned above our new method is to study algebraically isomonodromic extensions and to relate them to rigidity properties of the Hodge data on local systems. The notion of an isomonodromic extension of a bundle with a flat connection on a fibre $X_{s}$ of a good holomorphic map $f\colon X\to S$ to a relative flat bundle over a neighborhood of $s\in S$ is well-established in the theory of non-linear differential equations; its connection to Hodge theory has beed studied in particular in [LL24a].
@@
Let $f\colon X\to S$ be a good holomorphic map as above. Let $\mathbb{L}_{s}$ be a $K$-local system on $X_{s}$ with semi-simple algebraic monodromy group. Then the following conditions are equivalent:
@@
Recall that if $\mathbb{L}_{s}$ underlies a ()-$({\mathbb{Z}}){\mathbb{C}}\text{-}{\sf PVHS}$ then by [Del71, Sec. 4.2] its algebraic monodromy group is semi-simple, see Theorem 5.1(v). We do not know whether the lattice $\mathbb{L}_{{\mathcal{O}}_{K}}$ in ii) can be chosen as an extension of $\mathbb{L}_{s,{\mathcal{O}}_{K}}$.
@@
Let $S$ be a finite étale covering of the moduli stack of hyperbolic punctured compact Riemann surfaces with universal curve $f\colon X\to S$. Fix $s\in S$. Then a semi-simple local system $\mathbb{L}_{s}$ on $X_{s}$ such that the orbit $\pi_{1}(S,s)\cdot[\mathbb{L}_{s}]$ is finite, essentially extends to $X$.
@@
Acknowledgment. It is a pleasure to thank Alexander Petrov for numerous enlightening discussions and for sharing his insights into isomonodromic deformations of Hodge structures. In particular, he alerted us on a gap in a previous version of Appendix A. We thank Claude Sabbah for kindly helping us with references on variations of Hodge structure. We thank Daniel Litt for various helpful discussions on local systems in particular on the condition 5) in Theorem 1.1. We are grateful to Takuro Mochizuki for explaining to us what is known and what can be expected about families of pluri-harmonic metrics. We thank Xinwen Zhu for an enlightening discussion on the analogy between Theorem 1.1 and his own theorem on the preservation of the de Rham property of a $p$-adic local system, see [LZ17].
@@ 2.1. Group theory
We recall after Anderson [And74] some criteria when a finitely generated normal subgroup $H\subset G$ of a finitely generated abstract group $G$ induces an injective homomorphism $\widehat{H}\to\widehat{G}$ after profinite completion.
@@
()$(\star)\ \ \ \ \{1\}=\Gamma_{0}\subset\Gamma_{1}\subset\Gamma_{2}\subset\cdots\subset\Gamma_{s}=\Gamma$
@@
The first part is [And74, Prop. 4]. For the second part let us assume that $\Gamma$ is an extension of $s>0$ finitely generated free groups as above. We first assume that $s=1$ that is that $\Gamma$ is free. Then the surjection $G\to\Gamma$ has a section and we can conclude as before. For $s\geq 2$, we define $G_{s-1}$ as the inverse image in $G$ of $\Gamma_{s-1}$. By induction on $s$ we can assume that $\hat{H}\to\hat{G}_{s-1}$ is injective. By the case $s=1$, $\hat{G}_{s-1}\to\hat{G}_{s}$ is injective. Thus the composite $\hat{H}\to\hat{G}_{s-1}\to\hat{G}_{s}$ is injective as well. ∎
@@ 2.2. Geometric examples
Let $X$ and $S$ be connected complex manifolds and let $f\colon X\to S$ be a good holomorphic map as defined in Section 1. Fix a point $x\in X$ and set $s=f(x)$. By Thom’s first isotopy lemma [GM88, Thm. I.1.5] or Ehresmann’s theorem in the proper case, $f$ is a topological fibre bundle. We obtain an exact sequence of homotopy groups
@@
Let $\mathbb{L}$ be a local system on $S$ with finite fibres. If $S$ is an Artin $K\pi 1$ and $i>0$ is an integer, there exists a finite étale cover $S^{\prime}\to S$ such that the pull-back map
@@
By [And74, Prop. 18], the centers of $\pi_{1}(X_{s},x)$ and of $\widehat{\pi_{1}(X_{s},x)}$ are trivial. Thus by Proposition 2.3 it is sufficient to see that the homomorphism $\pi_{2}(S,s)\to\pi_{1}(X_{s},x)$ is trivial, which follows as by hyperbolic geometry one shows that $\pi_{1}(X_{s},s)$ does not contain any normal non-trivial abelian subgroup.
@@ 3.1. Integral Fixed Part Theorem
Let $X$ and $S$ be quasi-projective complex manifolds and let $f\colon X\to S$ be a projective holomorphic submersive map with connected fibres. Fix a point $\tilde{s}$ in the universal cover $\tilde{S}$ of $S$ and let $s\in S$ be its image. We say that a property essentially holds if the property holds after replacing $S$ by a finite quotient cover $S^{\prime}\to S$ of $\tilde{S}\to S$, $s$ by the image of $\tilde{s}$ in $S^{\prime}$ and $X$ by $X\times_{S}S^{\prime}$. We denote by $X_{s}$ the fibre over $s$.
@@
As $f$ is a topological fibre bundle, $\pi_{1}(S,s)$ acts on the cohomology $H^{i}(X_{s},{\mathbb{Q}})$. We denote by $\widehat{S}$ the formal completion of $S$ along $s$ and by $\widehat{X}$ the formal completion of $X$ along $X_{s}$.
@@
Let $f:X\to S$ be a proper good holomorphic map with $S$ as in I) in Section 1. Let $s$ be a complex point of $S$. Let $\xi\in H^{2i}(X_{s},{\mathbb{Q}})$ be a Hodge class on $X_{s}$. Then the following conditions are equivalent:
@@
    the Gauß-Manin flat deformation of $\xi$ in $H^{2i}_{dR}(\hat{X}/\hat{S})$ lies in the subbundle
    $F^{i}H^{2i}(\hat{X}/\hat{S}):=H^{2i}(\hat{X}/\hat{S},\Omega^{\geq i}_{\hat{X}/\hat{S}})\subset H^{2i}_{dR}(\hat{X}/\hat{S}).$
@@
We have to check that after replacing $S$ by a finite quotient covering $S^{\prime}$ of $\tilde{S}$ the class $\zeta$ becomes integral. The obstruction for integrality is the image of $\zeta$ in $H^{2i}(X,{\mathbb{Q}}/{\mathbb{Z}})$. There exists an $N>0$ such that this image is induced by an element of
@@
We now address 5) $\Rightarrow$ 1). Recall that the Hodge or Noether-Lefschetz locus $\mathrm{NL}$ is the closed analytic subset of the étale space associated to $R^{2i}f_{*}{\mathbb{Q}}$ consisting of Hodge cycles, see [CDK95, Intro.]. Restricted to any irreducible component of $\mathrm{NL}$ the morphism $\mathrm{NL}\to S$ is finite and unramified. The assumption 5) implies that there exists only one irreducible component $W$ of $\mathrm{NL}$ containing the given Hodge cycle $\xi$ and that $W\to S$ is étale at $\xi$. Then $W\to S$ is a finite, surjective, unramified holomorphic map of reduced, irreducible analytic spaces of the same dimension with a complex manifold as codomain, so it is a finite étale covering of complex manifolds. We conclude that the orbit $\pi_{1}(S,s)\cdot\xi$ consists of the finitely many elements of the fibre of $W\to S$ over $s$. See also [BEK14, Intro.] and references therein. ∎
@@ 4.1. Main group theoretic result
We denote by $[\rho_{H}]\in\mathrm{Rep}(H)$ the isomorphism class of a representation $\rho_{H}\colon H\to{\mathrm{GL}}(V)$. There is a natural action of $\Gamma$ on the set of isomorphism classes of representations $\mathrm{Rep}(H)$ by conjugation by lifts to $G$ of elements in $\Gamma$.
@@
Assume that the map of profinite completions $\widehat{H}\to\widehat{G}$ is injective. Let $\rho_{H}\colon H\to{\mathrm{GL}}(V)$ be a representation with semi-simple algebraic monodromy group. The following condition are equivalent.

### v5

@@ Abstract.
We formulate and prove a non-abelian analog of Deligne’s Fixed Part theorem on Hodge classes, revisiting previous work of Jost–Zuo, Katzarkov–Pantev and Landesman–Litt. To this aim we study algebraically isomonodromic extensions of local systems, and we relate them to variations of Hodge structures, for example we show that the Mumford-Tate group at a generic point stays constant in an algebraically isomonodromic extension of a variation of Hodge structure.
@@ 1. Introduction
On a quasi-projective complex manifold, a rational Betti class is called Hodge if it is integral and lies as a de Rham class in the right levels of the Hodge and the weight filtrations. Let $f\colon X\to S$ be a projective morphism between quasi-projective complex manifolds. Deligne’s Fixed Part Theorem asserts that a Hodge class on a fibre $X_{s}$ has finite orbit under the fundamental group of $S$ based at $s$ if and only if it extends to a Hodge class on $X$ after base change to a finite étale cover of $S$. Here we have to make the technical assumption that $S$ topologically has the properties of Artin’s elementary neighborhoods, see Theorem 3.1 for a precise formulation.
@@
If $R\subset K$ is a subring we denote by $(R)K\text{-}{\sf PVHS}$ those polarized variations of $K$-Hodge structure $(\mathbb{L},{\mathsf{F}},Q)$ such that the local system $\mathbb{L}$ can be defined over $R$ in the classical sense, that is that its monodromy representation can be defined over $R$.
Let $f:X\to S$ be an algebraic morphism between quasi-projective complex manifolds. We say that $f$ is a good algebraic morphism if there is a smooth extension $\overline{f}:\overline{X}\to S$ of $f$ such that $\overline{X}\supset X$ is quasi-projective smooth and $D=\overline{X}\setminus X$ is a relative normal crossings divisor. We prove our main results assuming that $f$ satisfies one of the following conditions, and we assume for the rest of the introduction that this is the case.
Ass: $f:X\to S$ is a good algebraic morphism and at least one of the following properties holds for $f$:
@@
Let $f:X\to S$ be a good algebraic morphism fulfilling Ass, $s$ be a point in $S$. Let $(\mathbb{L}_{s},{\mathsf{F}}_{s},Q_{s})$ be a $({\mathbb{Z}})K\text{-}{\sf PVHS}$ on the fiber $X_{s}$, where $K\in\{{\mathbb{Q}},{\mathbb{R}},{\mathbb{C}}\}$. Then the following conditions are equivalent:
@@
    the filtration ${\mathsf{F}}_{s}$ extends to a (relative) Griffiths transversal filtration by subbundles $\hat{\mathsf{F}}$ of the formal isomonodromic Deligne extension $)(\hat{\mathcal{E}},\hat{\nabla}\colon\hat{\mathcal{E}}\to\hat{\Omega}^{1}_{\hat{\overline{X}}/\hat{S}}(\log D)(\hat{\mathcal{E}}$)).
Here “essentially extends” means that it extends after base change with respect to a finite, étale, surjective covering of $S$, see Subsection 3.1 for a precise formulation. In 5) $\hat{S}$ is the formal scheme of $S$ along $s$ and $\hat{\overline{X}}$ the formal scheme of $\overline{X}$ along $\overline{X}_{s}$. The formal isomonodromic Deligne extension (^,^)($\hat{\mathcal{E}}$$\hat{\nabla}$) is the formal completion of the log flat bundle on $\overline{X}_{\Delta}$ associated by Deligne [Del70, Prop. 5.4, Rmk. 5.5] to the local system on $X_{\Delta}$ which canonically extends $\mathbb{L}_{s}$. Recall that the filtration ${\mathsf{F}}_{s}$ canonically extends to a filtration by subbundles of the Deligne extension over $\overline{X}_{s}$ by Schmid’s nilpotent orbit theorem, see [CK89, (2.1)], so part 5) can be interpreted as a formal deformation problem for the latter filtration of bundles from $\overline{X}_{s}$ to $\hat{\overline{X}}$.
@@
In order to show the implication 4) $\Rightarrow$ 1) we use Simpson’s non-abelian Hodge loci. In fact, he considers only the case $f$ projective, and we generalize in Appendix A his theory to the case of a good holomorphic map $f:X\to S$, that is $X$ and $S$ are complex manifolds, $f$ is submersive and admits a relative compactification $\overline{f}:\overline{X}\to S$ such that $\overline{f}$ is a projective morphism and $D=\bar{X}\setminus X$ is a relative normal crossings divisor. The analogous implication 5) $\Rightarrow$ 1) was suggested to us by D. Litt, see also [Lit24, Cor. 4.4.3].
@@
As mentioned above our new method is to study algebraically isomonodromic extensions and to relate them to rigidity properties of the Hodge data on local systems. The notion of an isomonodromic extension of a bundle with a flat connection on a fiber $X_{s}$ of a good holomorphic map $f\colon X\to S$ to a relative flat bundle over a neighborhood of $s\in S$ is well-established in the theory of non-linear differential equations; its connection to Hodge theory has been studied in particular in [LL24a].
@@
Let $f\colon X\to S$ be a good algebraic morphism fulfilling Ass. Let $\mathbb{L}_{s}$ be a $K$-local system on $X_{s}$ with semisimple algebraic monodromy group. Then the following conditions are equivalent:
@@
Recall that if $\mathbb{L}_{s}$ underlies a $({\mathbb{Z}}){\mathbb{C}}\text{-}{\sf PVHS}$ then by [Del71, Sec. 4.2] its algebraic monodromy group is semisimple, see Theorem 5.1(v). We do not know whether the lattice $\mathbb{L}_{{\mathcal{O}}_{K}}$ in ii) can be chosen as an extension of $\mathbb{L}_{s,{\mathcal{O}}_{K}}$.
@@
Let $S$ be a finite étale covering of the moduli stack of hyperbolic punctured compact Riemann surfaces with universal curve $f\colon X\to S$. Fix $s\in S$. Then a semisimple local system $\mathbb{L}_{s}$ on $X_{s}$ such that the orbit $\pi_{1}(S,s)\cdot[\mathbb{L}_{s}]$ is finite, essentially extends to $X$.
@@
Acknowledgment. It is a pleasure to thank Alexander Petrov for numerous enlightening discussions and for sharing his insights into isomonodromic deformations of Hodge structures. In particular, he alerted us on a gap in a previous version of Appendix A. We thank Claude Sabbah for kindly helping us with references on variations of Hodge structure. We thank Daniel Litt for various helpful discussions on local systems in particular on the condition 5) in Theorem 1.1. We are grateful to Takuro Mochizuki for explaining to us what is known and what can be expected about families of pluri-harmonic metrics. We thank Xinwen Zhu for an enlightening discussion on the analogy between Theorem 1.1 and his own theorem on the preservation of the de Rham property of a $p$-adic local system, see [LZ17]. It is a pleasure to acknowledge the influence of [LL24b] on our work. Finally, we thank the referee for a friendly, precise and helpful report.
@@ 2.1. Group theory
We recall after Anderson [And74] some criteria under which a finitely generated normal subgroup $H\subset G$ of a finitely generated abstract group $G$ induces an injective homomorphism $\widehat{H}\to\widehat{G}$ after profinite completion.
@@
$(\star)\ \ \ \ \{1\}=\Gamma_{0}\subset\Gamma_{1}\subset\Gamma_{2}\subset\cdots\subset\Gamma_{s}=\Gamma$
@@
The first part is [And74, Prop. 4]. For the second part let us assume that $\Gamma$ is an extension of $s>0$ finitely generated free groups as above. We first assume that $s=1$ that is that $\Gamma$ is free. Then the surjection $G\to\Gamma$ has a section and we can conclude as before. For $s\geq 2$, we define $G_{s-1}$ as the inverse image in $G$ of $\Gamma_{s-1}$. By induction on $s$ we can assume that $\hat{H}\to\hat{G}_{s-1}$ is injective. By the case $s=1$, $\hat{G}_{s-1}\to\hat{G}_{s}$ is injective. Thus, the composite $\hat{H}\to\hat{G}_{s-1}\to\hat{G}_{s}$ is injective as well. ∎
@@ 2.2. Geometric examples
Let $X$ and $S$ be connected complex manifolds and let $f\colon X\to S$ be a good holomorphic map as defined in Section 1. Fix a point $x\in X$ and set $s=f(x)$. By Thom’s first isotopy lemma [GM88, Thm. I.1.5] or Ehresmann’s theorem in the proper case, $f$ is a topological fiber bundle. We obtain an exact sequence of homotopy groups
@@
Let $\mathbb{L}$ be a local system on $S$ with finite fibers. If $S$ is an Artin $K\pi 1$ and $i>0$ is an integer, there exists a finite étale cover $S^{\prime}\to S$ such that the pull-back map
@@
By [And74, Prop. 18], the centers of $\pi_{1}(X_{s},x)$ and of $\widehat{\pi_{1}(X_{s},x)}$ are trivial. Thus, by Proposition 2.3 it is sufficient to see that the homomorphism $\pi_{2}(S,s)\to\pi_{1}(X_{s},x)$ is trivial, which follows as by hyperbolic geometry $\pi_{1}(X_{s},x)$ does not contain any normal non-trivial abelian subgroup. To see this classical fact, one may e.g. apply [Kat92, Thm. 2.3.6] and the triviality of the center of any finite index subgroup of $\pi_{1}(X_{s},x)$.
@@ 3.1. Integral Fixed Part Theorem
Let $X$ and $S$ be quasi-projective complex manifolds and let $f\colon X\to S$ be a projective holomorphic submersive map with connected fibers. Fix a point $\tilde{s}$ in the universal cover $\tilde{S}$ of $S$ and let $s\in S$ be its image. We say that a property essentially holds if the property holds after replacing $S$ by a finite cover $S^{\prime}\to S$ with $\tilde{S}\to S^{\prime}\to S$, $s$ by the image of $\tilde{s}$ in $S^{\prime}$ and $X$ by $X\times_{S}S^{\prime}$. We denote by $X_{s}$ the fiber over $s$.
@@
As $f$ is a topological fiber bundle, $\pi_{1}(S,s)$ acts on the cohomology $H^{i}(X_{s},{\mathbb{Q}})$. We denote by $\widehat{S}$ the formal completion of $S$ along $s$ and by $\widehat{X}$ the formal completion of $X$ along $X_{s}$.
@@
Let $f:X\to S$ be a projective, submersive morphism of quasi-projective complex manifolds with $S$ as in I) in Section 1. Let $s$ be a complex point of $S$. Let $\xi\in H^{2i}(X_{s},{\mathbb{Q}})$ be a Hodge class on $X_{s}$. Then the following conditions are equivalent:
@@
[…]
@@
We have to check that after replacing $S$ by a finite covering $S^{\prime}\to S$ with $\tilde{S}\to S^{\prime}\to S,$ the class $\zeta$ becomes integral. The obstruction for integrality is the image of $\zeta$ in $H^{2i}(X,{\mathbb{Q}}/{\mathbb{Z}})$. There exists an $N>0$ such that this image is induced by an element of
@@
We now address 5) $\Rightarrow$ 1). Recall that the Hodge or Noether-Lefschetz locus $\mathrm{NL}$ is the closed analytic subset of the étale space associated to $R^{2i}f_{*}{\mathbb{Q}}$ consisting of Hodge cycles, see [CDK95, Intro.]. Restricted to any irreducible component of $\mathrm{NL}$ the morphism $\mathrm{NL}\to S$ is finite and unramified. The assumption 5) implies that there exists only one irreducible component $W$ of $\mathrm{NL}$ containing the given Hodge cycle $\xi$ and that $W\to S$ is étale at $\xi$. Then $W\to S$ is a finite, surjective, unramified holomorphic map of reduced, irreducible analytic spaces of the same dimension with a complex manifold as codomain, so it is a finite étale covering of complex manifolds. We conclude that the orbit $\pi_{1}(S,s)\cdot\xi$ consists of the finitely many elements of the fiber of $W\to S$ over $s$. See also [BEK14, Intro.] and references therein. ∎
@@ 4.1. Main group theoretic result
We denote by $\rho_{H}]\in\mathrm{Rep}(H)$ the isomorphism class of a representation $\rho_{H}\colon H\to{\mathrm{GL}}(V)$. There is a natural action of $\Gamma$ on the set of isomorphism classes of representations $\mathrm{Rep}(H)$ by conjugation by lifts to $G$ of elements in $\Gamma$.
@@
Assume that the map of profinite completions $\widehat{H}\to\widehat{G}$ is injective. Let $\rho_{H}\colon H\to{\mathrm{GL}}(V)$ be a representation with semisimple algebraic monodromy group. The following conditions are equivalent.